so this is the algebra problem of the week #12
the problem is as follows
A man has a bag of coins. He meets three people.The first one gets 1/2 the coins +2. The second get 1/2 the coins +2. The third gets 1/2 the coins +2. When the man returns home he has two coins in his bag. Write and solve an equation to determine how many he started with.

Answer by Boreal(15235)

He began with 44 coins.
He started with x coins.

He gave away (x/2) +2, subtracted from x, leaves him with (x/2)-2
Second person gets half of that, which (x/4)-1+2=(x/4)+1
Now he is left with the difference, which is (x/4)-3
He gives away half of that (x/8)-3/2+2=(x/8)+(1/2)
He had (x/4)-3 and he gave away (x/8) +1/2. When you subtract, (x/4)-(x/8)=(x/8) And -3 -1/2=-(7/2)
He is left with (x/8)-(7/2)=2 because he has two coins left.
Multiply everything in the equation by 8 to clear fractions.
x-28=16 (7/2)*8=28
Add 28 to both sides.
x=44 coins.
Gives away half (22)+ 2 and has given away 24. He has 20 left.
He gives away half (10) +2 or 12, and he has 8 left.
He gives away half (4) plus 2 or 6, and he has 2 left.
Don't let the fractions of coins, or what appear to be fractional coins worry you. Just keep working it out and see what happens.